OpenStudy (anonymous):

/Find the complex fourth root of 100 - 100 sqrt 3 i

OpenStudy (anonymous):

write in trig form first

OpenStudy (anonymous):

What is the trig form?

OpenStudy (anonymous):

\(r\left(\cos(\theta)+i\sin(\theta)\right)\)

OpenStudy (anonymous):

\(r=\sqrt{a^2+b^2}=\sqrt{100^2+100^2}=100\sqrt2\)

OpenStudy (anonymous):

and in this case \(\theta=-\frac{\pi}{3}\) or if you are working in degree \(-60\)

OpenStudy (anonymous):

or maybe even \(300\) degrees

OpenStudy (anonymous):

ok, I have the trig form. From there what? The sqrt3 i?